OpenStudy (anonymous):

if 6 counters are 2/3 of a set how many counters are in one and two thirds set?

OpenStudy (mertsj):

6/2/3=x/5/3

OpenStudy (mertsj):

Now cross multiply: 6(5/3) = 2/3x 10=2/3x Multiply by 3/2 15=x

OpenStudy (anonymous):

15

OpenStudy (anonymous):

thank you soo much!